Patient 4 month old child who weighs12 lb On hand: Erythromycin oral s Ordered: Erythromycin...

  

  1. Patient 4 month old child who weighs12 lb
    On hand: Erythromycin oral s
    Ordered: Erythromycin suspension 65 mg PO q12huspension 100 mg/ 2.5 mL. According to the package insert the following are usual dosages for children over 3 months of age. For mild to moderate ear, nose throat infections, either 25 mg/kg/day in divided doses every 12 h, or 20mg/kg/dfay in divided doses every 8 h.

Is this order safe. YES_____________ NO____________ Why? Show results of calculations.

Amount to administer if safe:__________________
Answer #1

The weight of patient = 12 lb=5 kg

Recommended safe dose= 25 mg/kg/day in divided doses every 12 h, or 20mg/kg/day in divided doses every 8 h.

Thus the daily dose for the patient= 25X 5=125 mg or 20x 5= 100 mg per day in divided doses every 8 h

Ordered: Erythromycin suspension 65 mg PO q12h = 130 mg per day

This is not safe order as its more than the recommended safe dose, hence need to adjust dose as per the weight of patient within safe dose range.

Further for  130 mg per day the required ml will be=?

100 mg contains 2.5 mL

1 mg= 2.5/100

130mg will contain=(2.5/100) X130=3.25 ml